# Continuously Learning New Words in Automatic Speech Recognition

### v3

@@ I Introduction
In the last decade, ASR systems improved tremendously in terms of word error rate (WER) due to more data, more computing power, and better architectures [1, 2]. However, these systems are still far from perfect. Although in principal end-to-end systems are open-vocabulary systems, when using appropriate modeling units, such as byte-pair encoded (BPE) characters, in practice, words not seen during training are often not reliably recognized. Typical errors are observed in the categories of cross-lingual words (e.g. ’upgeloaded’ is mixing German and English), numbers (e.g. 1945 vs. 19:45 vs. \$19.45), acronyms (e.g. ICASSP), named entities and domain-specific special words (as they occur in specialized meetings or lectures). The word error rate (WER) is only slightly affected by these errors because they are rather infrequent. However, these words are important for understanding the content as they carry a lot of information that is lost when they are misrecognized. To measure this, we must not only evaluate WER but also recall, precision, and F1 score for these words, i.e., how often are they recognized if they occur and how often do they produce false positives.
@@ II Related Work
An approach to new-word learning of ASR systems is context biasing [3, 4, 5, 6]. The model is biased with a list of words. Depending on the approach, this list is used in different ways. The problem is that the scale of such approaches for continual learning is limited.
@@ VI Acknowledgements
This research was supported in part by a grant from Zoom Video Communications, Inc. The authors gratefully acknowledge the support. Furthermore, the projects on which this research is based were funded by the Federal Ministry of Education and Research (BMBF) of Germany under the numbers 01IS18040A (OML) and 01EF1803B (RELATER) and we want to thank Ngoc Quan Pham for providing his framework NMTGMinor to train our systems.

### v4

@@ I Introduction
In the last decade, ASR systems improved tremendously in terms of word error rate (WER) due to more data, more computing power, and better architectures [1, 2]. However, these systems are still far from perfect. Although in principal end-to-end systems are open-vocabulary systems, when using appropriate modeling units, such as byte-pair encoded (BPE) characters, in practice, words not seen during training are often not reliably recognized. Typical errors are observed in the categories of cross-lingual words (e.g. ’upgeloaded’ is mixing German and English), numbers (e.g. 1945 vs. 19:45 vs. \$19.45), acronyms (e.g. ICASSP), named entities and domain-specific special words (as they occur in specialized meetings or lectures). The word error rate (WER) is only slightly affected by these errors because they are rather infrequent. However, these words are important for understanding the content as they carry a lot of information that is lost when they are misrecognized. Furthermore, proper interpretation of these words is critical for downstream processing such as in speech translation [3, 4]. To measure this, we must not only evaluate WER but also recall, precision, and F1 score for these words, i.e., how often are they recognized if they occur and how often do they produce false positives.
@@ II Related Work
An approach to new-word learning of ASR systems is context biasing [5, 6, 7, 8, 9, 10]. The model is biased with a list of words. Depending on the approach, this list is used in different ways. The problem is that the scale of such approaches for continual learning is limited.
@@ VI Acknowledgements
This research was supported in part by a grant from Zoom Video Communications, Inc. Furthermore, the projects on which this research is based were funded by the Federal Ministry of Education and Research (BMBF) of Germany under the number 01EF1803B (RELATER), the Horizon research and innovation program of the European Union under grant agreement No 101135798 (Meetween), and the KIT Campus Transfer GmbH (KCT) staff in accordance with the collaboration with Carnegie – AI. The authors gratefully acknowledge the support. We thank Ngoc Quan Pham for providing his framework NMTGMinor to train our systems.